OpenStudy (anonymous):

what is the density of a rock with a mass of 2.5 grams, and a volume of 10 ml ? *Calculate the density of your rock in grams per cubic centimeter (1 cm^3 = 1 ml)*

OpenStudy (anonymous):

@shabazz \[\frac{ mass }{ volume }=\frac{ 2.5g }{ 10ml }=\frac{ .25g }{ ml }\times \frac{ 1ml }{ cm^3}=\frac{ .25 g}{ cm^3 }\]

OpenStudy (anonymous):

So the answer is .25 grams per cubic centimeter.